### v1

# arrows.meta, chains, positioning

Abstract
Reinforcement learning is well known for its ability to model sequential tasks and learn latent data patterns adaptively. Deep learning models have been widely explored and adopted in regression and classification tasks. However, deep learning has its limitations such as the assumption of equally spaced and ordered data, and the lack of ability to incorporate graph structure in terms of time-series prediction. Graphical neural network (GNN) has the ability to overcome these challenges and capture the temporal dependencies in time-series data. In this study, we propose a novel approach for predicting time-series data using GNN and monitoring with Reinforcement Learning (RL). GNNs are able to explicitly incorporate the graph structure of the data into the model, allowing them to capture temporal dependencies in a more natural way. This approach allows for more accurate predictions in complex temporal structures, such as those found in healthcare, traffic and weather forecasting. We also fine-tune our GraphRL model using a Bayesian optimisation technique to further improve performance. The proposed framework outperforms the baseline models in time-series forecasting and monitoring. The contributions of this study include the introduction of a novel GraphRL framework for time-series prediction and the demonstration of the effectiveness of GNNs in comparison to traditional deep learning models such as RNNs and LSTMs. Overall, this study demonstrates the potential of GraphRL in providing accurate and efficient predictions in dynamic RL environments.
Index Terms:
Graph Neural Networks, Reinforcement Learning, Intelligent Monitoring, Bayesian Optimization
I Introduction
Machine learning (ML) applications in healthcare drive towards automation of clinician tasks, enhance their capabilities and access to patient care [1]. The research community widely adopts advanced ML mechanisms such as Federated Learning due to their ability to maintain data privacy while building robust state-of-the-art server models [2]. Another ML mechanism Reinforcement Learning(RL) has shown great success in enhancing the prediction performance [3] and can assist in sequential decision-making tasks [4]. One of the most important applications of RL in dynamic environments is in the control of autonomous systems, such as robots and drones. These systems must be able to navigate and interact with their environment in real time, and RL can be used to train them to make optimal decisions based on sensor data. For example, RL can be used to train a robot to navigate through a cluttered environment, or a drone to fly in a safe and efficient manner.
In multiple domains, such as healthcare, traffic, and weather forecasting, early warning systems (EWS) are crucial to detect and diagnose issues early before they become more severe and challenging to handle. Real-time monitoring data is analyzed, and alerts are issued when certain thresholds are crossed, indicating a potential issue. Self-learning systems, like RL-based EWS, are particularly useful as they can adapt to changes in a condition over time and continually refine their predictions. By integrating an RL-based EWS into clinical decision-making, clinicians can make more informed decisions and take proactive measures to prevent adverse health events from occurring. RL-based EWS have been successful in predicting hospital readmissions, sepsis detection, and adverse drug events. In traffic, EWS can alert drivers to congestion, accidents, or other hazards, allowing them to plan alternative routes and avoid delays. In weather forecasting, EWS can provide advance notice of severe weather conditions, allowing individuals and communities to prepare and take necessary precautions to minimize damage and loss of life. Ultimately, EWS are vital in various domains to enable proactive measures, prevent damage, and enhance safety and quality of life.
RL environment with observation space defining states at each time step can be analogised to time-series data. Modelling time-series data for monitoring and predicting future states has been widely explored by deep learning models. Deep learning models, such as Recurrent Neural Networks (RNNs) and Long Short-Term Memory (LSTM) networks [5], play a significant role in predicting time-series data. These models are able to capture the temporal dependencies between time steps in the data, allowing them to make accurate predictions. The RNNs and LSTM architectures are designed to handle sequential data, making them well-suited for time-series prediction tasks.
Static temporal modelling, which is widely used in deep learning models like RNNs and LSTM networks, can help forecast and monitor potential issues and detect early warnings in various domains, including healthcare, traffic, and weather forecasting. These models can capture the temporal dependencies between time steps in the data and make accurate predictions. However, RNNs and LSTMs have some limitations when it comes to time-series prediction tasks, such as the assumption of equally spaced and ordered data and the lack of ability to incorporate graph structure where the data has a complex temporal structure.
This study proposes a novel framework called GraphRL, which combines reinforcement learning (RL) with graph-based neural networks to forecast the future states of a dynamic environment. Specifically, the framework uses a type of graph-based neural network called Temporal Graphical Convolutional Network (T-GCN) to capture the temporal dependencies between different time steps in a more natural way, by explicitly incorporating the graph structure of the data. One of the advantages of T-GCN is that it can also incorporate other forms of information, such as node and edge attributes, which can be useful for time-series prediction tasks where there is additional information associated with each time step. This is particularly important for deep learning challenges in Q-learning function approximation, where neural networks are used to approximate the Q-function. To evaluate the effectiveness of GraphRL, its performance is compared to other state-of-the-art models on a range of benchmark datasets. The results show that GraphRL outperforms the baseline models on most of the datasets, demonstrating the potential of using graph-based neural networks in RL. The contributions of this study are:
  • A novel and generic GraphRL framework with predictive and monitoring capabilities for early warnings in a complex environment.
  • A customised and generic RL environment to predict future states of a dynamic environment such as healthcare, traffic, and weather forecast.
  • A new paradigm of virtual monitoring of predicted states in RL, allowing for timely interventions by communicating with the appropriate team.
The rest of the paper is organised as Section II presents the existing works on self-learning systems and prediction tasks. Section III presents the research problem. In Section IV, the proposed GraphRL framework and its algorithm are discussed. Section V presents the datasets used to evaluate the generic GraphRL framework and the baseline models are presented. Predictive RL environment and agent performances are presented and compared with the baseline models in Section VI-A. The proposed framework hyperparameters are fine-tuned with Bayesian optimisation in Section VII. Finally, the paper concludes in Section VIII.
II Related Works
II-A Self Learning Systems
[6] proposed a framework representing a user as an RL agent and monitoring their behaviours in mobile health. The authors deployed two agents: a user agent to model the user decision-making process and an app agent to assist the user in achieving their goals by maximising the frequency of the user performing the behaviour. The proposed framework extracted different intervention strategies based on its self-learning experience. [7] adopted RL to model maladaptive eating behaviours for application-delivered mindfulness training intervention. [8] proposed Model Implants for Rapid Reflective Other-agent Reasoning/Learning (MIRROR) using RL to bootstrap human modelling. The MIRROR framework can quickly learn human behaviour from demonstrations. In personalising human behaviour, [9] discussed CalFit, a fitness app, built on the RL algorithm to generate personalised daily step goals by monitoring changes in behaviour and implementing dynamic goal setting and self-monitoring. Li et al. [10] propose a method that combines imitation learning with reinforcement learning to improve the success rate of autonomous driving in urban scenarios, which are more complex than simple scenarios such as highways and parking lots. The method decomposes the action space into low-level and high-level action spaces and uses a DQN-based reinforcement learning approach to execute specific imitation learning policies. The approach also introduces a new reward for high-level action selection.
II-B Early Detection of Patient Deterioration
Early detection of vital signs deterioration is key to the timely invention and avoiding clinical deterioration in acutely ill patients in hospitals. Traditional patient monitoring is to report individual vital signs of patients which states their current clinical status. For example, vital signs such as temperature, pulse, respiratory rate, and mean arterial pressure are considered continuous predictors for emergency department patients [11]. New patient monitoring algorithms analyze multiple features from physiological signals. They can produce an index that serves as a predictive or prognostic measure for a specific critical health event or physiological instability [12, 13]. Ashrafian et al. presented a case series where wireless remote vital signs monitoring systems on medical and surgical wards could reduce the time to detect deteriorating patients [14]. Kellett et al. [15] presented numerous studies as evidence of the deterioration of patients is often missed even though it is usually preceded by worsening vital signs. Although vital signs are important to measure there remain, challenges as well as opportunities for new technology like wearable devices [15, 16]. Early recognition of patient deterioration in hospitals is crucial in reducing morbidity and preventing long-term disability. Early clinical prediction is typically based on manually calculated screening metrics that weigh these parameters, such as Early warning scores (EWS) [17]. EWS is considered one of the metrics to estimate early detection. There have been reports published that EWS are excellent predictors of cardiac arrest [18], ICU transfer [19], death in ICU [20] and as well as 30-day mortality and length of stay in ICU [21].
II-C Vital Signs Prediction
[22] proposed two prediction models for mortality prediction and length of ICU stay respectively. The authors applied six traditional machine learning models to classify survived or not in mortality prediction, and ICU length of stay prediction with two machine learning models. Similarly, [23] proposed five predictive models in five different scenarios to predict vital signs such as heart rate, blood pressure, respiration rate and oxygen saturation of covid-19 patients in ICU. [24] proposed the patient’s clinical time-series variable prediction framework with multi-head transformers. The authors discussed four different scenarios: clinical time series forecasting, in-hospital mortality prediction, 24-hour mortality prediction and Length Of Stay prediction. [25] proposed a deep learning approach, DeepVS, for radio frequency-based vital signs sensing. The author combined 1D CNN and attention models to exploit local features and temporal correlations to estimate heart and respiration rates.
The use of self-learning algorithms in sequential decision-making tasks has garnered attention in the research community as they aid healthcare practitioners in making personalised interventions and prescribing appropriate medication. While deep learning methods such as CNN, RNN, and transformers have shown success in vital signs prediction, they have limitations such as assuming equally spaced and ordered data, and not being able to incorporate graph structure. Therefore, the aim of this study is to introduce an adaptive self-learning prediction framework that utilises the power of GNNs [26, 27] to capture temporal dependencies between timesteps and enable real-time monitoring with reinforcement learning.
III Research Problem
The research problem is to overcome the deep learning challenges and predict future states of a complex and dynamic RL environment and learn the latent behaviour patterns of data adaptively.
Definition 1
Let $V$ be set of $n$ vital parameters $V_{t}=\{v^{1},v^{2},...,v^{n}\}$ at time $t$ determining the health status of a subject $S$. The vital parameters are continuous time-series data changing dynamically and inferring the subject’s health status. For time-series forecasting, the continuous vital parameters are split into time windows such as $T=\{V_{t-2},V_{t-1},V_{t},V_{t+1},V_{t+2},.....,V_{t+n}\}$. Non-linear models are trained with the historic data in time windows $\{V_{t-2},V_{t-1},V_{t}\}$ to predict future $\{V_{t+1},V_{t+2},.....,V_{t+n}\}$.
Definition 2
Based on the training, the subject $S$ has a set of learning agents following the Markov Decision Process (MDP) to monitor and learn the patterns of their vital parameters $V_{t}=\{v^{1},v^{2},...,v^{n}\}$ continuously. This can be defined as a 5-tuple $M$= ($S,A,P,R,\gamma$) constituting $S$ as finite state space, in which $s_{t}\epsilon S$ denotes the state of an agent at time $t$, $A$ denotes set of actions defined for the agent in which $a_{t}\epsilon A$ denotes the action taken by the agent at time $t$, $P$ is a Markovian transition function which denotes agent transits from state $s$ to state $s^{\prime}$ while performing an action $a$ through $P(s,a,s^{\prime}):S\times A\times S$. The transition function $P$ leads to a reward function $R:S\times A\rightarrow\Re$ which returns an immediate reward $R(s,a)$ for the action $a$ taken in a state $s$ defined in Equation 1.${\gamma}$ is a discount factor that focuses on immediate rewards instead of future rewards, which remains between 0 and 1.
${\displaystyle R(s_{t},a_{t})=\sum_{t=0}^{\infty}\gamma^{t}r_{t},}$ (1)
returns an immediate reward R(s, a) for the action taken in a state s defined in Equation 1
Figure 1: GraphRL Framework
IV GraphRL Framework
To address the research problem, a novel graphical neural network (GNN) enabled reinforcement learning (RL) framework is proposed. In the graph-enabled RL framework(GraphRL), two GNNs are deployed: one for forecasting time-series data and another for Q-function approximation as shown in Fig. 1. The proposed framework is demonstrated in Fig. 1 in which the interaction between an environment and an AI agent is illustrated. As discussed in the research problem, finite MDP is adopted to formulate the process of modelling current and past states.
IV-A Predictive GraphRL Environment
The primary objective of the proposed study is to learn from the past and current states of a dynamic environment and predict the future states of the complex environment. To achieve this objective, we propose a predictive monitoring environment which is responsible for defining the observation space with state $s_{t}^{i}\epsilon S$ where $i=0,1,2,...n$, action space with actions $a_{t}^{j}\epsilon A$ where $j=1,2,3,...m$, and rewards $R$ for each action taken by the agent as it transitions from a state $s_{t}$ to $s_{t+1}$ in a real-world scenario. For example, consider a subject in a dynamic environment whose current state is denoted as $V_{t}={v^{1},v^{2},...,v^{n}}$ at time $t$. Similarly, the subject holds historical data of their state at times $t-1,t-2,t-3,...,t-n$. In traditional reinforcement learning formulations, the monitoring environment is a static entity that cannot forecast future states, which might affect the subjects in the environment.
IV-A1 T-GCN Forecast
Forecasting the future states of a subject before a few time steps can revolutionise the most dynamic industries such as gaming, healthcare, and so on by identifying the deteriorating state of the subject in the environment. To predict the future state in a reinforcement learning environment, a temporal graph convolutional network(T-GCN) is adopted. The graphical network is trained with past and current states at their timestamps in a supervised approach as shown in Equation 3. The training process also includes the features leading to those states.
$y=f(b+\sum_{i=1}^{n}v_{i}.w_{i})$ (2)
$\begin{split}y(v)=\sum_{i=1}^{n}Activation1(b+w_{i}v_{i})\\ y(v)=Activation1(\frac{e^{y_{i}}}{\sum_{j=1}^{k}e^{y_{j}}})\end{split}$ (3)
IV-A2 Static Spatial-temporal modelling
A two-layered graphical network is adopted for Spatial-temporal modelling, a spatial modelling layer is based on a graphical convolutional network (GCN), and a temporal layer based on recurrent neural networks (RNN) is configured. The spatial layer is responsible to capture spatial features among nodes which are input states $s_{t}^{i}\epsilon S$. This can be achieved by constructing Fourier transform filter and it acts on the graph nodes and its first-order neighbourhood. In this study, a static graph temporal signal is adopted in which the node positions in the graph remain the same and the label information is dynamic. The spatial layer is to set the static graph with nodes as input states $s_{t}^{i}\epsilon S$. The two-layered GCN model is defined in Equation 4.
$f(X,A)=\sigma(\hat{A}Relu(\tilde{A}XW_{0})W_{1})$ (4)
Where X is the input matrix, A represents the graph adjacency matrix, $\hat{A}$ and $\tilde{A}$ represent the preprocessing step and self-connection structure respectively. $W_{0},W_{1}$ represents weights of the first and second layers of ST-GCN, and $\sigma(\cdot)$. $Relu()$ is an activation function.
Temporal modelling is based on the RNN variant gated recurrent unit (GRU) [28] which has a simple structure and faster training ability. In the GRU model, an update gate $z_{t}$ controls the degree of information retrieved from the previous state and a reset gate $r_{t}$ controls the degree of ignoring the status information at the previous moment are configured as shown in Equation 5.
$z_{t}$ $=\sigma_{g}(W_{z}[f(X_{t},A),h_{t-1}]+b_{z})$ (5)
$r_{t}$ $=\sigma_{g}(W_{r}[f(X_{t},A),h_{t-1}]+b_{r})$
${\hat{h}}_{t}$ $=\phi_{h}(W_{h}[f(X_{t},A),h_{t-1}](r_{t}\odot h_{t-1})+b_{h})$
$h_{t}$ $=z_{t}\odot h_{t-1}+(1-z_{t})\odot{\hat{h}}_{t}$
${\displaystyle x_{t}}:$ input vector, ${\displaystyle h_{t}}:$ output vector, ${\displaystyle{\hat{h}}_{t}}:$ candidate activation vector, ${\displaystyle z_{t}}:$ update gate vector, ${\displaystyle r_{t}}:$ reset gate vector, ${\displaystyle W},{\displaystyle U},{\displaystyle b}:$ parameter matrices and vector, $\sigma_{g}:$ The original is a sigmoid function, ${\displaystyle\phi_{h}}$ The original is a hyperbolic tangent.
In the training process, the T-GCN model predicts the future states at time $t+1,t+2,t+3,...,t+n$ as $\hat{Y_{t}}$ and compares with the real-data ${Y_{t}}$. This determines the loss function of the graph network [29] as shown in Equation 6. To avoid over-fitting problems in the training process, the loss function is optimised with L2 regularisation $L_{reg}$ and a hyperparameter $\lambda$.
$loss=||\hat{Y_{t}}-{Y_{t}}||+\lambda L_{reg}$ (6)
IV-B Algorithm I
Algorithm 1 Predictive GraphRL Environment
1:Input: time series data $\mathcal{D}=\{s_{t-n},..,s_{t-2},s_{t-1},s_{t},\}$; a set of labels $\mathcal{K}=\{1,2,\dots,K\}$
2:Output: Predicted time series data of $\mathcal{K}$, a set of labels, in the form of states $\{s_{t+1},s_{t+2},s_{t+3},s_{t+4}\}$.
3:Define $forecast\_model\leftarrow T-GCNModel$
4:$Train(forecast\_model)\leftarrow forecast\_model(D)$
5:$\{s_{t+1},s_{t+2},s_{t+3},s_{t+4}\}\leftarrow forecast\_model(predict)$
6:$Initialization:\ observation\_space={s_{t}^{i}\epsilon S},action\_space={a_{t}\epsilon A},rewardR$
7:$Set\ monitor\_length=N$
8:if $actionisappropriate$ then
9:     $R\leftarrow+reward$
10:else
11:     $R\leftarrow-reward$
12:end if
13:$monitor\_length\leftarrow N-1$
14:$s_{t+1}\leftarrow s_{t}(monitor\_length)$
15:if $N=0$ then
16:     $done=True$
17:else
18:     $done=False$
19:end if
20:$visualize(a_{t},R,vital\ signs)$
21:$initial\_state\leftarrow s_{t}[0]$triangleright reset environment
The proposed GraphRL environment is defined in Algorithm 1, which configures the observation space, action space, and reward policy based on predicted states. The T-GCN model is defined in lines 1-3 to train and predict time series forecasting data. The class is initialized and boundaries for observation space, action space, rewards, and monitoring length are set in lines 4-5. Lines 6-10 describe the reward policy for learning agent actions and how agents are rewarded positively or negatively. Finally, lines 11-19 address monitoring length, visualization of agent performance, and resetting the environment if necessary.
IV-C GraphRL Agent
In this study, the game learning agent Deep Q-Networks (DQN) algorithm is used. The algorithm was introduced by Google’s DeepMind for playing Atari game to play games by just observing the screen without any training or prior knowledge about those games. In this algorithm, the Q-Learning functions’ approximation is computed using the proposed T-GCN model and the learning agent gets rewarded based on the graph network prediction of the right action for the current state.
IV-C1 Q-Function Approximation
T-GCN model used in this study to approximate the Q-Function for each action in the action space as shown in Fig. 1. The model is configured with parameters such as the relu activation function, mean square error as loss function, and Adam optimiser. The model gets trained with the state and its corresponding action. The learning agent performs an action $a_{t}\epsilon A$ for a transition from state $s_{t}$ to $s_{t}^{{}^{\prime}}$ and achieves a reward $R$ for the action. In this transition process, the maximum of the Q-function in Equation 7 is calculated, and the discount of the calculated value uses a discount factor $\gamma$ to suppress future rewards and focus on immediate rewards. The discounted future reward is added to the current reward to get the target value. The difference between the current prediction from the neural networks and the calculated target value provides a loss function. The loss function is a deviation of the predicted value from the target value and it can be estimated from Equation 8. The square of the loss function allows for the punishment of the agent for a large loss value.
$Q^{\pi}(s,a)=E_{\pi}\biggl{\{}\ \sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t},\pi(s_{t}))|s_{0}=s,a_{0}=a\biggl{\}}$ (7)
${loss=(\underbrace{R+\gamma\cdot max(Q^{\pi^{*}}(s,a))}_{target\_value}-\underbrace{Q^{\pi}(s,a)}_{predicted\_value})^{2}}$ (8)
IV-C2 Exploration and Exploitation
the concepts of exploration and exploitation are at odds with each other. Exploration involves randomly selecting actions that have not been performed before to uncover more possibilities and enhance the agent’s understanding. Exploitation, on the other hand, entails selecting actions based on past experiences and knowledge to maximize rewards. To balance the trade-off between exploration and exploitation, different strategies such as the greedy algorithm, epsilon-greedy algorithm, optimistic initialization, and decaying epsilon-greedy algorithm are employed. This study proposes controlling the exploration rate by multiplying the decay by the exploration rate. This approach reduces the number of explorations during execution as the agent learns patterns and maximizes rewards to achieve high scores. As the T-GCN model is retrained with previous experiences in the replay, the decay is multiplied by the exploration rate depending on the agent’s ability to predict the right actions. All these parameters are defined as hyper-parameters for DQN learning agents.
IV-C3 GraphRL Agent Algorithm
Algorithm 2 Learning Agent
1:Initialize $\gamma,\epsilon,\epsilon_{decay},\epsilon_{min},memory=\emptyset,batch\_size$
2:Define $model\leftarrow T-GCN\_model$
3:$memory\leftarrow append(s_{t},a_{t},R,s_{t+1})$
4:if $np.random.rand<\epsilon$ then $\triangleright$ Exploration
5:     $action\_value\leftarrow random(a_{t})$
6:elsenormal-$\triangleright$ Exploitation
7:     $action\_value\leftarrow model.predict(s_{t})$
8:end if
9:$minibatch\leftarrow random(memory,batch\_size)$
10:for $s_{t},a_{t},R,s_{t+1},done$ in $minibatch$ do
11:     $target\leftarrow R$
12:     if not done then
13:         $target\leftarrow R+\gamma(max(model.predict(s_{t+1})))$
14:     end if
15:     $target\_f\leftarrow model.predict(s_{t})$
16:     $target\_f[a_{t}]\leftarrow target$
17:     $model.fit(s_{t},target\_f)$
18:end for
19:if $\epsilon$geq $\epsilon_{min}$ then
20:     $\epsilon$= $\epsilon_{decay}$
21:end if
The Learning Agent algorithm 2 encapsulates methods for function approximation, memorization, replay, exploration, and exploitation. In lines 1-2, hyperparameters required for the agent and T-GCN model for Q-function approximation are initialized. Memorization and replay, explained in line 3, involve storing T-GCN experience, including state, action, reward, and next_state, for later retraining of the model using the replay method. The algorithm (lines 4-8) predicts an action using either exploration or exploitation methods. Finally, lines 9-21 retrieve a batch of previous experiences from memory, process and retrain the neural network model based on the earlier defined hyperparameters.
IV-C4 Implementation Algorithm
Algorithm 3 Proposed GraphRL Framework Implementation
1:Input:
2:$\mathcal{C}={1,2,\dots,C}$: set of subjects
3:$\mathcal{V}={1,2,\dots,V}$: set of vital signs
4:$\mathcal{M}={1,2,\dots,M}$: number of episodes
5:Output: Rewards achieved by Agents in each episode.
6:$env\leftarrow ForecastingEnvironment()$triangleright Algorithm 1
7:$agent\leftarrow LearningAgent()$triangleright Algorithm 2
8:for episode $m\in\mathcal{M}$ do
9:     $state\leftarrow env.reset()$
10:     $score=0$
11:     for time in range(timesteps) do
12:         $a_{t}\leftarrow agent.action(s_{t})$
13:         $s_{t+1},R,done\leftarrow env.step(a_{t})$
14:         $agent.memorize(s_{t},a_{t},R,s_{t+1})$
15:         $s_{t}\leftarrow s_{t+1}$
16:         if done then
17:              $print(m,score)$
18:              $break;$
19:         end if
20:     end for
21:     $agent.replay(batch\_size)$
22:end for
The Algorithm  3 corresponds to the previous two algorithms,  1 and  2, and implements the proposed GraphRL monitoring framework. It takes as input a set of subjects with their vital signs and the number of training episodes for the agents. The output of the algorithm is the cumulative reward score achieved by the learning agent in each episode. The first two lines of the algorithm create instances of the Predictive GraphRL environment and GraphRL agent. Lines 3-17 consist of nested for loops with conditional statements to determine if an episode is complete. The outer loop iterates through each episode, resetting the environment to its initial state and the score to zero. The inner loop iterates through timesteps, which represent the current state’s time, and calls the methods defined in Algorithms  1 and  2 to predict the current state’s action, reward the agent for its predicted action, retrieve the next state, and store previous experiences. Finally, the replay method is called to update the neural network model with the stored experiences.
IV-D Bayesian Optimisation
Bayesian optimisation is a global optimisation method that uses a probabilistic model to guide the search for the optimal solution. The model is updated as new data points are sampled and evaluated, allowing the algorithm to improve its predictions over time by fine-tuning the $L_{reg}$, $\lambda$ and minimising the loss function defined in Equation 6. The basic idea behind Bayesian optimisation is to model the objective function, f(x), as a Gaussian process (GP). The GP model is used to predict the objective function value at any point x, given the observations of the function at other points. The prediction is given by the posterior distribution of the GP, which is a Gaussian distribution with mean and variance given by Equation 9.
$\mu(x)$ $=k(x,X)^{T}(K+\sigma^{2}I)^{-1}y$ (9)
$\sigma^{2}(x)$ $=k(x,x)-k(x,X)^{T}(K+\sigma^{2}I)^{-1}k(X,x)$
Where X is the matrix of previously sampled points, y is the vector of corresponding function values, K is the Gram matrix of the covariance function evaluated at X, and $\sigma^{2}$ is the noise level in the function evaluations. The next point to sample is chosen based on an acquisition function, which balances the trade-off between exploration and exploitation. Common acquisition functions include the probability of improvement and the expected improvement. Given a set of observed points (X,y) and a Gaussian process prior, Bayesian optimisation seeks the point x* that minimises the loss function value, given by Equation 10. The optimisation process continues iteratively, sampling new points and updating the GP model until a stopping criterion is met.
${\displaystyle EI(x)=\mathbf{E}[max(0,f(x)-f(x^{*}))]}$ (10)
Where x* is the current best point.
V Experiment
The primary objective of this study is to overcome deep learning challenges such as the assumption of equally spaced and ordered data [30] and the lack of ability to incorporate graph structure where the data has a complex temporal structure [31]. These challenges are particularly relevant in domains such as health, weather, and traffic where it is important to analyze temporal patterns and make accurate forecasts for early warning systems. However, traditional deep learning models often fail to capture these complex patterns, limiting their effectiveness in these critical domains.
The proposed GraphRL framework is evaluated on three different forecasting applications: heart rate prediction, traffic forecast, and weather forecast. The framework predicts future events in the form of states and optimizes actions based on those predictions. The observation space is customized for each application and actions for the agent are pre-defined. The agent receives a reward for correctly predicting a state and communicating with the relevant team. The proposed approach is a generic framework that can be applied to monitor and predict time-series data and train an RL agent to learn the latent patterns of the monitoring process. The baseline models for comparison include traditional deep learning models such as GRU, LSTM, and RNNs.
V-A Datasets
In this study, the proposed framework is evaluated with three publicly available datasets related to healthcare, traffic and weather.
  • Wearable Stress and Affect Detection (WESAD) dataset [32]: This is a publicly available dataset that contains physiological and self-report data collected from 17 participants during various stress and affective states.The dataset contains data collected from various wearable sensors, including a chest-mounted electrocardiogram (ECG) sensor, a wrist-mounted photoplethysmogram (PPG) sensor, and an accelerometer.
  • Los Angeles (LA) traffic dataset [33]: This is a publicly available dataset that contains data on traffic flow in the Los Angeles metropolitan area. The dataset was collected by the Los Angeles Department of Transportation (LADOT) and contains information on traffic counts, traffic speeds, and travel times for a variety of purposes, including traffic prediction, traffic flow analysis, and transportation planning.
  • Large-Scale Traffic and Weather Events (LSTW) dataset [34]: It contains information on traffic conditions and extreme weather events in the United States. In terms of weather events, the dataset has several types including rain, snow, storm, cold weather event, etc.
V-B Baseline Models
  • In Li et al.’s work [35], the authors suggested the ELMA method for multi-agent activity forecasting through graph neural networks that incorporate spatiotemporal data using simultaneous energy-based learning.
  • Ma et al. [36] proposed a technique for predicting multi-agent driving behaviour in different scenarios based on self-supervised domain knowledge.
  • In Jiang et al.’s research [37], they conducted a study on internet traffic prediction with distributed multi-agent learning and utilized both LSTM and gated recurrent unit (GRU). Their GRU-based distributed multi-agent learning scheme outperformed LSTM.
V-C Evaluation Metrics
Mean Absolute Error (MAE) is a commonly used regression metric that determines the average magnitude of errors between the predicted and actual values in a given dataset. It is calculated by taking the average of the absolute differences between the predicted and actual values, which results in a single value. Another popular regression metric is the Root Mean Squared Error (RMSE), which measures the average magnitude of the differences between the predicted and actual values. RMSE is calculated as the square root of the mean of the squared differences between the predicted and actual values. Mean Absolute Percentage Error (MAPE) is another regression metric that calculates the average absolute percentage error between the predicted and actual values. It is calculated as the average of the absolute differences between the predicted and actual values, expressed as a percentage of the actual values. Cumulative Rewards, on the other hand, is a performance metric used in reinforcement learning that quantifies the total rewards obtained by an agent within a specific time frame or number of actions. It is computed by summing up all the rewards that the agent received during the given period or a number of actions. The experiments conducted in this study employed Python version 3.7.6 and utilized several packages, including TensorFlow, Keras, Open AI Gym, and stable_baselines3.
TABLE I: Performance of the proposed framework in health forecasting
15 Min 30 Min 45 Min 60 Min
ELMA [35] MAE 6.2 6.2 6.2 6.13
MAPE 13.91 13.91 13.91 13.91
RMSE 8.75 8.75 8.75 8.67
GRU [36] MAE 0.95 0.95 0.97 0.98
MAPE 5.47 5.48 5.51 5.5
RMSE 1.25 1.25 1.27 1.28
GRU-Based Multi-Agent [37] MAE 1.02 1.02 1.25 1.65
MAPE 8 3.47 4.53 5.27
RMSE 2.46 2.58 2.69 3.09
GraphRL (Ours) MAE 0.56 0.87 0.68 0.7
MAPE 2.8 2.9 2.65 3.98
RMSE 1.18 1.47 1.3 1.32
VI Results and Analysis
In this section, the proposed framework performance in terms of time series forecasting and RL monitoring is compared to the baseline models in each application.
VI-A Predictive GraphRL Performance
Healthcare Forecasting: The proposed framework is evaluated to monitor health status by predicting future vital signs such as heart rate. Based on the sensor data and other clinical parameters such as ECG, Respiration, the time series prediction of the heart rate is conducted. The predicted values of heart for the next one hour are break-down into different time intervals (15 minutes, 30 minutes, 45 minutes, 60 minutes). Each of these time interval values acts as an observation for the GraphRL agent to monitor and communicate with the appropriate emergency team. The observation space of the vital sign, action space of different emergency teams and rewards for the agent actions in the predictive GraphRL environment are defined based on the modified early warning scores (MEWS) [38]. For the evaluation process, the WESAD dataset is adopted to conduct time series forecasting of heart rate. The proposed T-GCN in the predictive GraphRL environment performs better than the other baseline frameworks ELMA, GRU, and GRU-Based Multi-Agent as shown in Tab. I. It achieves the lowest MAE, MAPE and RMSE values in all the time intervals.
TABLE II: Performance of the proposed framework in traffic forecasting
15 Min 30 Min 45 Min 60 Min
ELMA [35] MAE 6.73 6.73 6.73 6.72
MAPE 6.73 15.14 15.14 15.07
RMSE 6.72 9.4 9.4 9.39
GRU [36] MAE 1.04 1.04 1.04 1.04
MAPE 6.04 6.01 5.96 6.1
RMSE 1.36 1.36 1.36 1.36
GRU-Based Multi-Agent [37] MAE 1.85 1.85 1.96 1.82
MAPE 6.07 5.7 4.93 6.07
RMSE 2.88 3.21 3.43 3.54
GraphRL (Ours) MAE 0.65 0.78 0.64 0.8
MAPE 4.1 7.85 5.65 7.99
RMSE 1.27 1.22 1.26 1.41
Traffic Forecasting: The goal of the proposed framework is to predict traffic using the predictive GraphRL environment. The system takes in data with the following features: EventId, Type, Severity, TMC, Description, StartTime, EndTime, TimeZone, LocationLat, LocationLng, Distance, AirportCode, Number, Street, Side, City, County, State, and ZipCode. The observation space includes the current traffic state, which is represented by the traffic events and their severity in a particular region. The actions referred to possible traffic management strategies, such as altering traffic light timings or changing the speed limit. The rewards are defined based on the efficiency of the chosen strategy, such as reduced travel time or decreased congestion. For all the baseline models and the proposed framework, the MAE, MAPE, and RMSE values are reported for forecasting at 15, 30, 45, and 60-minute intervals. As shown in Tab. II, T-GCN outperforms the other models for all the forecasting intervals with the lowest MAE, MAPE, and RMSE values. The second-best performer is the GRU-Based Multi-Agent model, followed by GRU and ELMA.
TABLE III: Performance of the proposed framework in weather forecasting
15 Min 30 Min 45 Min 60 Min
ELMA [35] MAE 6.69 6.69 6.69 6.65
MAPE 6.69 15.02 15.02 14.99
RMSE 6.69 9.39 9.39 9.34
GRU [36] MAE 1.03 1.03 1.04 1.04
MAPE 5.96 5.94 5.93 6.02
RMSE 1.36 1.35 1.36 1.36
GRU-Based Multi-Agent [37] MAE 1.65 1.65 1.85 2.02
MAPE 7.32 4.71 4.89 5.86
RMSE 2.76 2.99 3.16 3.43
GraphRL (Ours) MAE 0.61 0.83 0.66 0.75
MAPE 3.95 5.88 5.15 7.99
RMSE 1.23 1.12 1.28 1.26
Weather Forecasting: In weather forecasting, the goal of the proposed framework is to use past weather data to predict future weather events and to optimise actions based on those predictions. In the predictive environment, the observation space is configured based on both the traffic and weather events datasets, including the event type, severity, start time, end time, location (latitude and longitude), and timezone. The actions represent the decisions the RL agent can take based on the observation space. For example, the agent could decide to issue a warning or alert for severe weather, adjust traffic signals or road signs, or change the speed limit on certain roads. The agent could receive a reward for correctly predicting severe weather and issuing a timely warning. Using the proposed GraphRL framework allows modelling the relationships between different weather events and their impact on traffic in a more efficient way than traditional machine learning methods. The GraphRL agent learns from these relationships to make better decisions and improve its predictions over time. Comparing the different models, T-GCN had the best performance across all metrics and different time intervals: 15, 30, 45, and 60 minutes, followed by GRU-Based Multi-Agent, GRU, and ELMA. The results show that the T-GCN model had the lowest MAE, MAPE, and RMSE values for all forecasting intervals, indicating its superior forecasting performance compared to the other models as shown in Tab. III.
VI-B GraphRL Agent Performance
TABLE IV: Proposed GraphRL Performance
AI Agents WESAD
LAM Traffic
Forecasting
US Weather
Forecasting
Q Learning 43130 28840 39480
PPO 39480 33945 29480
A2C 41195 22845 40615
Double DQN 42615 25600 33945
DDPG 44600 34590 39945
DQN 41986 35219 40985
GraphRL 48790 36195 53145
The proposed RL agent was enabled with T-GCN and its performance is compared with other traditional RL agents as shown in Tab. IV. The table provides a comparison of different AI agents and their performance on three different datasets: WESAD, LAM Traffic Forecasting, and US Weather Forecasting. The performance of each agent is measured by a score, which is the total score achieved by the agent on the task over ten episodes. From the table, it can be seen that the proposed GraphRL agent is the most efficient agent on the WESAD dataset, as it scored the highest score. The DDPG and Q-Learning agents have the second-highest score on the WESAD dataset. On the LAM Traffic Forecasting dataset, the Q-Learning agent scored the lowest, and the proposed GraphRL agent scored the highest. On the US Weather Forecasting dataset, the A2C agent scored the lowest, while the GraphRL agent scored the highest. The GraphRL agent has outperformed other RL agents in all three predictive and monitoring applications.
[ xlabel=Episodes, ylabel=Score, legend pos=south east, legend style=font=, yticklabel style=/pgf/number format/fixed, ymin=-20000, ymax=65000, width=height=6cm, xtick=1,2,3,4,5,6,7,8,9,10, xticklabels=1,2,3,4,5,6,7,8,9,10]
[color=blue, mark=*] coordinates (1, 38925) (2, 43595) (3, 46875) (4, 57280) (5, 32245) (6, 485) (7, 48035) (8, 48790) (9, 42885) (10, 54470) ;
[color=red, mark=square*] coordinates (1, -8810) (2, -8100) (3, -11845) (4, 31960) (5, 46295) (6, 32560) (7, 53640) (8, 36195) (9, 36480) (10, 46425) ;
[color=green, mark=triangle*] coordinates (1, 23780) (2, -6765) (3, 33480) (4, 43635) (5, 24925) (6, 2670) (7, 8990) (8, 53145) (9, 58530) (10, 41405) ;
Health, Traffic, Weather
Figure 2: GraphRL Agent Rewards Distribution
The performance of the GraphRL agent is measured by the episode score, which appears to be the total score achieved by the agent after a certain number of episodes. The breakdown of the proposed agent’s score in each episode of the three applications is presented and compared in Fig. 2. The agent’s performance on the WESAD dataset is relatively consistent, with the scores fluctuating between 32245 and 57280. On the LAM Traffic Forecasting dataset, the agent’s performance is relatively inconsistent, with the scores fluctuating between -11845 and 46295. On the US Weather Forecasting dataset, the agent’s performance is also relatively inconsistent, with the scores fluctuating between -6765 and 58530. This inconsistency of the scores is due to the exploration rate where the algorithm tries exploring all the actions randomly instead of using T-GCN model predictions.
VII Bayesian Optimisation Results
[ xlabel=Episode, ylabel=Rewards, legend pos=north west, legend style=font=, legend cell align=left, ymajorgrids=true, grid style=dashed, ymin=-15000, ymax=60000, scaled y ticks = false, y tick label style=/pgf/number format/sci, width=height=6cm]
[…]
Figure 3: Bayesian optimisation of $\alpha$ for GraphRL Agent
[ xlabel=Episode, ylabel=Rewards, legend pos=north west, legend style=font=, legend cell align=left, ymajorgrids=true, grid style=dashed, ymin=-15000, ymax=60000, scaled y ticks = false, y tick label style=/pgf/number format/sci, width=height=6cm]
[color=red,mark=x] coordinates (1, -944) (2, -667) (3, 919) (4, 4009) (5, 7860) (6, 14303) (7, 21405) (8, 28799) (9, 38095) (10, 48823) ;
[color=blue,mark=x] coordinates (1, -588) (2, 2080) (3, 3122) (4, 6573) (5, 9759) (6, 16085) (7, 25136) (8, 34720) (9, 46093) (10, 57388) ;
[color=green,mark=x] coordinates (1, 909) (2, -258) (3, 1398) (4, 5207) (5, 10434) (6, 17728) (7, 25584) (8, 34898) (9, 32468) (10, 42637) ;
[color=orange,mark=x] coordinates (1, -84) (2, 272) (3, 2355) (4, 5792) (5, 11162) (6, 17760) (7, 23002) (8, 31702) (9, 40771) (10, 52126) ;
[…]
Figure 4: Bayesian optimisation of $\gamma$ for GraphRL Agent
The results of Bayesian optimisation for the proposed GraphRL agent using different values of the learning rate parameter, $\alpha$, during different episodes are shown in Fig. 3. The values in the y-axis of the line chart represent the scores or rewards obtained by the agent during each episode. It can be observed that the performance of the agent varies for different values of alpha. For example, in episode 1, the agent performs better with alpha = 0.001 (20630) compared to the other values. Similarly, in episode 10, the agent performs better with alpha = 0.001 (97380) compared to the other values. These results suggest that the optimal value of alpha for the agent is $\alpha=0.001$, and Bayesian optimisation can be used to find the best value of alpha for a given task.
These results in Fig. 4 show the performance of an RL agent using temporal GCN for Q function approximation, using different values of the discount factor gamma. As we can see, the performance varies greatly depending on the value of the gamma chosen. A high value of gamma (0.95) results in poor performance, while lower values (0.75) result in better performance. This suggests that a lower discount factor is more appropriate, as it gives more weight to immediate rewards and less to future rewards. It also suggests that there is an optimal value of gamma, which would need to be further explored through more extensive experimentation.
VIII Conclusion
In conclusion, this study proposed a novel GraphRL framework based on a Temporal Graphical Convolutional Network (T-GCN) to forecast the future states of a dynamic RL environment and overcome the limitations of deep learning models in time-series prediction tasks. The proposed framework was fine-tuned with the Bayesian optimisation technique and was applied to three datasets: WESAD, LAM Traffic Forecasting, and US Weather Forecasting. The results showed that the proposed GraphRL agent outperformed other traditional RL agents in all three predictive and monitoring applications. The agent’s performance on the WESAD dataset was relatively consistent, while on the LAM Traffic Forecasting and US Weather Forecasting datasets, the agent’s performance was relatively inconsistent. The inconsistency of the scores was due to the exploration rate. The Bayesian optimisation process was able to fine-tune the $L_{reg}$, $\lambda$ and minimise the loss function to improve the performance of the GraphRL agent. The proposed GraphRL framework can be applied to a wide range of applications and can be further enhanced by incorporating other forms of information such as node and edge attributes for more accurate predictions.
Ethical Statement
There are no ethical issues.

### v2

@@
-arrows.meta, chains, positioning
+Graph-enabled Reinforcement Learning for Time Series Forecasting with Adaptive Intelligence
 Abstract
@@
 I Introduction
-Machine learning (ML) applications in healthcare drive towards automation of clinician tasks, enhance their capabilities and access to patient care [1].
-The research community widely adopts advanced ML mechanisms such as Federated Learning due to their ability to maintain data privacy while building robust state-of-the-art server models [2].
-Another ML mechanism Reinforcement Learning(RL) has shown great success in enhancing the prediction performance [3] and can assist in sequential decision-making tasks [4].
-One of the most important applications of RL in dynamic environments is in the control of autonomous systems, such as robots and drones.
-These systems must be able to navigate and interact with their environment in real time, and RL can be used to train them to make optimal decisions based on sensor data.
-For example, RL can be used to train a robot to navigate through a cluttered environment, or a drone to fly in a safe and efficient manner.
-In multiple domains, such as healthcare, traffic, and weather forecasting, early warning systems (EWS) are crucial to detect and diagnose issues early before they become more severe and challenging to handle.
-Real-time monitoring data is analyzed, and alerts are issued when certain thresholds are crossed, indicating a potential issue.
-Self-learning systems, like RL-based EWS, are particularly useful as they can adapt to changes in a condition over time and continually refine their predictions.
-By integrating an RL-based EWS into clinical decision-making, clinicians can make more informed decisions and take proactive measures to prevent adverse health events from occurring.
-RL-based EWS have been successful in predicting hospital readmissions, sepsis detection, and adverse drug events.
-In traffic, EWS can alert drivers to congestion, accidents, or other hazards, allowing them to plan alternative routes and avoid delays.
-In weather forecasting, EWS can provide advance notice of severe weather conditions, allowing individuals and communities to prepare and take necessary precautions to minimize damage and loss of life.
-Ultimately, EWS are vital in various domains to enable proactive measures, prevent damage, and enhance safety and quality of life.
-RL environment with observation space defining states at each time step can be analogised to time-series data.
-Modelling time-series data for monitoring and predicting future states has been widely explored by deep learning models.
-Deep learning models, such as Recurrent Neural Networks (RNNs) and Long Short-Term Memory (LSTM) networks [5], play a significant role in predicting time-series data.
-These models are able to capture the temporal dependencies between time steps in the data, allowing them to make accurate predictions.
-The RNNs and LSTM architectures are designed to handle sequential data, making them well-suited for time-series prediction tasks.
-Static temporal modelling, which is widely used in deep learning models like RNNs and LSTM networks, can help forecast and monitor potential issues and detect early warnings in various domains, including healthcare, traffic, and weather forecasting.
-These models can capture the temporal dependencies between time steps in the data and make accurate predictions.
-However, RNNs and LSTMs have some limitations when it comes to time-series prediction tasks, such as the assumption of equally spaced and ordered data and the lack of ability to incorporate graph structure where the data has a complex temporal structure.
-This study proposes a novel framework called GraphRL, which combines reinforcement learning (RL) with graph-based neural networks to forecast the future states of a dynamic environment.
-Specifically, the framework uses a type of graph-based neural network called Temporal Graphical Convolutional Network (T-GCN) to capture the temporal dependencies between different time steps in a more natural way, by explicitly incorporating the graph structure of the data.
-One of the advantages of T-GCN is that it can also incorporate other forms of information, such as node and edge attributes, which can be useful for time-series prediction tasks where there is additional information associated with each time step.
-This is particularly important for deep learning challenges in Q-learning function approximation, where neural networks are used to approximate the Q-function.
-To evaluate the effectiveness of GraphRL, its performance is compared to other state-of-the-art models on a range of benchmark datasets.
-The results show that GraphRL outperforms the baseline models on most of the datasets, demonstrating the potential of using graph-based neural networks in RL.
-The contributions of this study are:
-  • A novel and generic GraphRL framework with predictive and monitoring capabilities for early warnings in a complex environment.
-  • A customised and generic RL environment to predict future states of a dynamic environment such as healthcare, traffic, and weather forecast.
-  • A new paradigm of virtual monitoring of predicted states in RL, allowing for timely interventions by communicating with the appropriate team.
-The rest of the paper is organised as Section II presents the existing works on self-learning systems and prediction tasks.
-Section III presents the research problem.
-In Section IV, the proposed GraphRL framework and its algorithm are discussed.
-Section V presents the datasets used to evaluate the generic GraphRL framework and the baseline models are presented.
-Predictive RL environment and agent performances are presented and compared with the baseline models in Section VI-A.
-The proposed framework hyperparameters are fine-tuned with Bayesian optimisation in Section VII.
-Finally, the paper concludes in Section VIII.
+The emergence of Machine Learning (ML) in healthcare signifies a paradigm shift towards automating clinician tasks and augmenting patient care capabilities [1].
+Amidst the evolving ML landscape, Federated Learning has gained traction for preserving data privacy while constructing sophisticated server models [2].
+Reinforcement Learning (RL), another ML strategy, has demonstrated substantial improvements in prediction performance and decision-making tasks [3, 4].
+RL’s application is particularly noteworthy in controlling autonomous systems, such as robots and drones, training them to make optimal decisions in real-time based on environmental sensor data.
+In various sectors, including healthcare, traffic, and weather forecasting, Early Warning Systems (EWS) play a pivotal role.
+They analyze real-time monitoring data and issue alerts for potential issues, facilitating proactive responses.
+RL-based EWS can adapt over time, refining their predictions and supporting clinical decision-making.
+This adaptability has proven effective in applications like predicting hospital readmissions and sepsis detection.
+Figure 1: Graphical Abstract
+Time-series data modeling, vital in monitoring and predicting future states, has seen advancements with deep learning models like Recurrent Neural Networks (RNNs) and Long Short-Term Memory (LSTM) networks [5].
+These models are adept at capturing temporal dependencies, yet face limitations in handling irregularly structured data and integrating complex graph structures.
+This study introduces the GraphRL framework, an innovative amalgamation of RL and Temporal Graphical Convolutional Networks (T-GCN), aiming to surpass the constraints of traditional deep learning models in time-series prediction.
+GraphRL’s design facilitates handling complex temporal structures and incorporates additional information such as node and edge attributes, as depicted in Fig.
+1.
+The GraphRL framework’s core contributions include:
+  • A versatile framework capable of providing early warnings and monitoring in complex settings.
+  • A customizable RL environment designed for effective forecasting in dynamic domains like healthcare and traffic systems.
+  • A novel approach to virtual monitoring of predicted states in RL, enhancing decision-making and intervention capabilities.
+Our comparative analysis with state-of-the-art models across various datasets showcases GraphRL’s superior performance, underscoring its potential as a versatile solution for time-series prediction challenges.
+The paper is organized as follows: Section II reviews existing literature on self-learning systems and prediction tasks.
+Section III outlines the research problem.
+The proposed GraphRL framework and its algorithm are detailed in Section IV.
+Section V describes the datasets and baseline models used for evaluation.
+The performance of the predictive RL environment and agent is compared with baseline models in Section VI.
+Section VII discusses the fine-tuning of the framework’s hyperparameters using Bayesian optimization.
+The paper concludes in Section VIII.
 II Related Works
 II-A Self Learning Systems
-[6] proposed a framework representing a user as an RL agent and monitoring their behaviours in mobile health.
-The authors deployed two agents: a user agent to model the user decision-making process and an app agent to assist the user in achieving their goals by maximising the frequency of the user performing the behaviour.
-The proposed framework extracted different intervention strategies based on its self-learning experience.
-[7] adopted RL to model maladaptive eating behaviours for application-delivered mindfulness training intervention.
-[8] proposed Model Implants for Rapid Reflective Other-agent Reasoning/Learning (MIRROR) using RL to bootstrap human modelling.
-The MIRROR framework can quickly learn human behaviour from demonstrations.
-In personalising human behaviour, [9] discussed CalFit, a fitness app, built on the RL algorithm to generate personalised daily step goals by monitoring changes in behaviour and implementing dynamic goal setting and self-monitoring.
-Li et al.
-[10] propose a method that combines imitation learning with reinforcement learning to improve the success rate of autonomous driving in urban scenarios, which are more complex than simple scenarios such as highways and parking lots.
-The method decomposes the action space into low-level and high-level action spaces and uses a DQN-based reinforcement learning approach to execute specific imitation learning policies.
-The approach also introduces a new reward for high-level action selection.
+Self-learning systems, particularly those utilizing Reinforcement Learning (RL), have seen significant advancements in various applications.
+For instance, Shin et al.
+[6] introduced a dual-agent framework in mobile health, effectively demonstrating user modeling and behavior intervention strategies.
+This work underscores the potential of RL in personalizing user experiences, a concept that aligns with our GraphRL framework’s goal of adaptive learning in dynamic environments.
+Similarly, Taylor et al.
+[7] applied RL in modeling maladaptive eating behaviors, further showcasing RL’s versatility in behavior prediction and modification.
+Chen et al.
+[8] developed the MIRROR framework, emphasizing the rapid learning capabilities of RL in human behavior modeling.
+These advancements set a precedent for our work in complex sequential decision-making tasks.
+Zhou et al.’s [9] CalFit app and Li et al.’s [10] method in autonomous driving highlight RL’s efficacy in personalized goal setting and complex urban scenario navigation, respectively, which are foundational to our GraphRL framework’s approach in handling dynamic and intricate patterns in data.
 II-B Early Detection of Patient Deterioration
-Early detection of vital signs deterioration is key to the timely invention and avoiding clinical deterioration in acutely ill patients in hospitals.
-Traditional patient monitoring is to report individual vital signs of patients which states their current clinical status.
-For example, vital signs such as temperature, pulse, respiratory rate, and mean arterial pressure are considered continuous predictors for emergency department patients [11].
-New patient monitoring algorithms analyze multiple features from physiological signals.
-They can produce an index that serves as a predictive or prognostic measure for a specific critical health event or physiological instability [12, 13].
-Ashrafian et al.
-presented a case series where wireless remote vital signs monitoring systems on medical and surgical wards could reduce the time to detect deteriorating patients [14].
-Kellett et al.
-[15] presented numerous studies as evidence of the deterioration of patients is often missed even though it is usually preceded by worsening vital signs.
-Although vital signs are important to measure there remain, challenges as well as opportunities for new technology like wearable devices [15, 16].
-Early recognition of patient deterioration in hospitals is crucial in reducing morbidity and preventing long-term disability.
-Early clinical prediction is typically based on manually calculated screening metrics that weigh these parameters, such as Early warning scores (EWS) [17].
-EWS is considered one of the metrics to estimate early detection.
-There have been reports published that EWS are excellent predictors of cardiac arrest [18], ICU transfer [19], death in ICU [20] and as well as 30-day mortality and length of stay in ICU [21].
+In the healthcare domain, early detection of patient deterioration is vital.
+Traditional vital signs monitoring, as discussed by Asiimwe et al.
+[11] and evaluated by Scully et al.
+[12], Baig et al.
+[13], and others, has laid the groundwork for our study.
+These works highlight the importance of continuous monitoring and early warning systems (EWS), which are integral to GraphRL’s objective.
+The limitations in existing methods, such as the need for manual calculations and the inability to handle large, unstructured data effectively, are addressed in our framework through the integration of GNNs, which can process complex temporal data more efficiently.
 II-C Vital Signs Prediction
-[22] proposed two prediction models for mortality prediction and length of ICU stay respectively.
-The authors applied six traditional machine learning models to classify survived or not in mortality prediction, and ICU length of stay prediction with two machine learning models.
-Similarly, [23] proposed five predictive models in five different scenarios to predict vital signs such as heart rate, blood pressure, respiration rate and oxygen saturation of covid-19 patients in ICU.
-[24] proposed the patient’s clinical time-series variable prediction framework with multi-head transformers.
-The authors discussed four different scenarios: clinical time series forecasting, in-hospital mortality prediction, 24-hour mortality prediction and Length Of Stay prediction.
-[25] proposed a deep learning approach, DeepVS, for radio frequency-based vital signs sensing.
-The author combined 1D CNN and attention models to exploit local features and temporal correlations to estimate heart and respiration rates.
-The use of self-learning algorithms in sequential decision-making tasks has garnered attention in the research community as they aid healthcare practitioners in making personalised interventions and prescribing appropriate medication.
-While deep learning methods such as CNN, RNN, and transformers have shown success in vital signs prediction, they have limitations such as assuming equally spaced and ordered data, and not being able to incorporate graph structure.
-Therefore, the aim of this study is to introduce an adaptive self-learning prediction framework that utilises the power of GNNs [26, 27] to capture temporal dependencies between timesteps and enable real-time monitoring with reinforcement learning.
+Vital signs prediction has been explored through various machine learning models.
+Alghatani et al.
+[14] and Youssef et al.
+[15] demonstrated the use of traditional machine learning in mortality prediction and vital signs forecasting, respectively.
+Harerimana et al.’s [16] work with multi-head transformers and Xie et al.’s [17] DeepVS model highlight the potential of deep learning in this domain.
+However, these methods often assume equally spaced and ordered data and lack the ability to incorporate complex graph structures, limitations our GraphRL framework aims to overcome.
+II-D Temporal Graphical Convolutional Networks (T-GCN)
+The integration of T-GCN within our GraphRL framework is pivotal.
+T-GCNs, known for their ability to capture temporal dependencies and complex relationships in graph-structured data, offer significant enhancements in processing time-series data [18].
+This technology addresses limitations in traditional deep learning models by effectively managing irregular time intervals and integrating additional contextual information (such as node and edge attributes) for richer data representation [19].
+The inclusion of T-GCN in GraphRL allows for a more nuanced understanding and prediction of dynamic systems [20], making it highly suitable for applications in healthcare monitoring, traffic forecasting, and weather prediction.
+The capability of T-GCNs to handle non-linear and complex temporal patterns [21] aligns with the core objectives of GraphRL, pushing the boundaries of current self-learning systems in real-world scenarios.
+In summary, while existing works in self-learning systems, patient deterioration detection, and vital signs prediction have laid a strong foundation, our GraphRL framework aims to address their limitations by introducing a novel approach that combines the strengths of GNNs and RL.
+This approach allows for a more sophisticated handling of temporal dependencies and real-time monitoring, which is crucial in dynamic environments such as healthcare, traffic management, and weather forecasting.
+The motivation behind the use of RL in our framework primarily arises from the need to tackle the challenges of multi-step time series prediction, where traditional supervised learning approaches may encounter limitations.
+Although supervised learning methods like GNN+Bert and GNN+TCN are indeed common and effective for time series forecasting, RL offers a unique advantage in dealing with situations where errors can accumulate over time, especially in dynamic environments.
+RL enables our predictive GraphRL Environment not only to forecast future states but also to actively influence decision-making, a capability particularly valuable in applications such as healthcare monitoring and the gaming industry.
 III Research Problem
-The research problem is to overcome the deep learning challenges and predict future states of a complex and dynamic RL environment and learn the latent behaviour patterns of data adaptively.
-Definition 1
-Let $V$ be set of $n$ vital parameters $V_{t}=\{v^{1},v^{2},...,v^{n}\}$ at time $t$ determining the health status of a subject $S$.
-The vital parameters are continuous time-series data changing dynamically and inferring the subject’s health status.
-For time-series forecasting, the continuous vital parameters are split into time windows such as $T=\{V_{t-2},V_{t-1},V_{t},V_{t+1},V_{t+2},.....,V_{t+n}\}$.
-Non-linear models are trained with the historic data in time windows $\{V_{t-2},V_{t-1},V_{t}\}$ to predict future $\{V_{t+1},V_{t+2},.....,V_{t+n}\}$.
-Definition 2
-Based on the training, the subject $S$ has a set of learning agents following the Markov Decision Process (MDP) to monitor and learn the patterns of their vital parameters $V_{t}=\{v^{1},v^{2},...,v^{n}\}$ continuously.
-This can be defined as a 5-tuple $M$= ($S,A,P,R,\gamma$) constituting $S$ as finite state space, in which $s_{t}\epsilon S$ denotes the state of an agent at time $t$, $A$ denotes set of actions defined for the agent in which $a_{t}\epsilon A$ denotes the action taken by the agent at time $t$, $P$ is a Markovian transition function which denotes agent transits from state $s$ to state $s^{\prime}$ while performing an action $a$ through $P(s,a,s^{\prime}):S\times A\times S$.
-The transition function $P$ leads to a reward function $R:S\times A\rightarrow\Re$ which returns an immediate reward $R(s,a)$ for the action $a$ taken in a state $s$ defined in Equation 1.${\gamma}$ is a discount factor that focuses on immediate rewards instead of future rewards, which remains between 0 and 1.
-${\displaystyle R(s_{t},a_{t})=\sum_{t=0}^{\infty}\gamma^{t}r_{t},}$ (1)
-returns an immediate reward R(s, a) for the action taken in a state s defined in Equation 1
+The research problem addresses deep learning challenges in predicting future states of a complex and dynamic Reinforcement Learning (RL) environment and adaptively learning latent behavior patterns of data.
+Definition 1 (Vital Parameters and Time-Series Forecasting)
+In the context of our framework, we consider a set $V$ of $n$ vital parameters, denoted as $V_{t}=\{v^{1},v^{2},\ldots,v^{n}\}$, which represent continuous time-series data reflecting the health status of a subject $S$.
+These vital parameters are dynamic and change over time, providing valuable insights into the subject’s well-being.
+To facilitate time-series forecasting, we segment these continuous vital parameters into time windows, denoted as $T$, which encompasses data points from the past ($V_{t-2},V_{t-1},V_{t}$) and extends into the future ($V_{t+1},V_{t+2},\ldots,V_{t+n}$).
+Non-linear models are trained on historical data within the time windows $\{V_{t-2},V_{t-1},V_{t}\}$ to predict future values $\{V_{t+1},V_{t+2},\ldots,V_{t+n}\}$.
+Definition 2 (Learning Agents and Markov Decision Process)
+Following the training phase, subject $S$ is associated with a group of learning agents that operate based on the principles of the Markov Decision Process (MDP).
+This MDP is a 5-tuple denoted as $M=(S,A,P,R,\gamma)$, and it forms the foundation for continuous monitoring and pattern learning of the vital parameters $V_{t}=\{v^{1},v^{2},\ldots,v^{n}\}$.
+Here’s a breakdown of each component:
+  • $S$ represents a finite state space, where $s_{t}\in S$ signifies the state of an agent at a specific time $t$,
+  • $A$ is the set of actions available to each agent, and $a_{t}\in A$ represents the action taken by the agent at time $t$,
+  • $P$ is a Markovian transition function $P(s,a,s^{\prime})$ that quantifies the probability of the agent transitioning from state $s$ to state $s^{\prime}$ while executing action $a$,
+  • $R$ is a reward function $R:S\times A\rightarrow\mathbb{R}$ that provides an immediate reward $R(s,a)$ for the action $a$ performed in state $s$,
+  • $\gamma$ is a discount factor, ranging between 0 and 1, which emphasizes immediate rewards over future rewards.
+$R(s_{t},a_{t})=\sum_{t=0}^{\infty}\gamma^{t}r_{t},$ (1)
+This equation returns the immediate reward $R(s,a)$ for the action taken in state $s$, as defined in Eq. 1.
 Figure 2: GraphRL Framework
@@
 In the graph-enabled RL framework(GraphRL), two GNNs are deployed: one for forecasting time-series data and another for Q-function approximation as shown in Fig.
-1.
+2.
 The proposed framework is demonstrated in Fig.
@@
 To predict the future state in a reinforcement learning environment, a temporal graph convolutional network(T-GCN) is adopted.
-The graphical network is trained with past and current states at their timestamps in a supervised approach as shown in Equation 3.
+The graphical network is trained with past and current states at their timestamps in a supervised approach as shown in Eq.
+3.
 The training process also includes the features leading to those states.
@@
 $\begin{split}y(v)=\sum_{i=1}^{n}Activation1(b+w_{i}v_{i})\\ y(v)=Activation1(\frac{e^{y_{i}}}{\sum_{j=1}^{k}e^{y_{j}}})\end{split}$ (3)
+Eq.
+2 describes a basic neural network neuron, with $y$ as the output, $f$ as the activation function, $b$ as the bias, $v_{i}$ as the input features, and $w_{i}$ as the weights; while Eq.
+3 involves an ’Activation1’ function, computing a weighted sum of inputs and normalizing these outputs, possibly into a probability distribution, akin to a softmax function.
 IV-A2 Static Spatial-temporal modelling
@@
 The spatial layer is to set the static graph with nodes as input states $s_{t}^{i}\epsilon S$.
-The two-layered GCN model is defined in Equation 4.
+The two-layered GCN model is defined in Eq.
+4.
 $f(X,A)=\sigma(\hat{A}Relu(\tilde{A}XW_{0})W_{1})$ (4)
@@
 Temporal modelling is based on the RNN variant gated recurrent unit (GRU) [22] which has a simple structure and faster training ability.
-In the GRU model, an update gate $z_{t}$ controls the degree of information retrieved from the previous state and a reset gate $r_{t}$ controls the degree of ignoring the status information at the previous moment are configured as shown in Equation 5.
+In the GRU model, an update gate $z_{t}$ controls the degree of information retrieved from the previous state and a reset gate $r_{t}$ controls the degree of ignoring the status information at the previous moment are configured as shown in Eq.
+5.
 $z_{t}$ $=\sigma_{g}(W_{z}[f(X_{t},A),h_{t-1}]+b_{z})$ (5)
@@
 In the training process, the T-GCN model predicts the future states at time $t+1,t+2,t+3,...,t+n$ as $\hat{Y_{t}}$ and compares with the real-data ${Y_{t}}$.
-This determines the loss function of the graph network [29] as shown in Equation 6.
+This determines the loss function of the graph network [23] as shown in Eq.
+6.
 To avoid over-fitting problems in the training process, the loss function is optimised with L2 regularisation $L_{reg}$ and a hyperparameter $\lambda$.
 $loss=||\hat{Y_{t}}-{Y_{t}}||+\lambda L_{reg}$ (6)
-IV-B Algorithm I
+IV-B Predictive GraphRL Environment Algorithm
 Algorithm 1 Predictive GraphRL Environment
-1:Input: time series data $\mathcal{D}=\{s_{t-n},..,s_{t-2},s_{t-1},s_{t},\}$; a set of labels $\mathcal{K}=\{1,2,\dots,K\}$
-2:Output: Predicted time series data of $\mathcal{K}$, a set of labels, in the form of states $\{s_{t+1},s_{t+2},s_{t+3},s_{t+4}\}$.
+1:Input: time series data $\mathcal{D}=\{s_{t-n},\ldots,s_{t-2},s_{t-1},s_{t}\}$; a set of labels $\mathcal{K}=\{1,2,\dots,K\}$
+2:Output: Predicted time series data of $\mathcal{K}$, a set of labels, in the form of states $\{s_{t+1},s_{t+2},s_{t+3},s_{t+4}\}$
 3:Define $forecast\_model\leftarrow T-GCNModel$
-4:$Train(forecast\_model)\leftarrow forecast\_model(D)$
+4:$Train(forecast\_model)\leftarrow forecast\_model(\mathcal{D})$
 5:$\{s_{t+1},s_{t+2},s_{t+3},s_{t+4}\}\leftarrow forecast\_model(predict)$
-6:$Initialization:\ observation\_space={s_{t}^{i}\epsilon S},action\_space={a_{t}\epsilon A},rewardR$
+6:$Initialization:observation\_space=\{s_{t}^{i}\in S\},action\_space=\{a_{t}\in A\},rewardR$
 7:$Set\ monitor\_length=N$
@@
 15:if $N=0$ then
-16:     $done=True$
+16:    $done\leftarrow True$
 17:else
-18:     $done=False$
+18:    $done\leftarrow False$
 19:end if
@@
 21:$initial\_state\leftarrow s_{t}[0]$triangleright reset environment
-The proposed GraphRL environment is defined in Algorithm 1, which configures the observation space, action space, and reward policy based on predicted states.
-The T-GCN model is defined in lines 1-3 to train and predict time series forecasting data.
-The class is initialized and boundaries for observation space, action space, rewards, and monitoring length are set in lines 4-5.
-Lines 6-10 describe the reward policy for learning agent actions and how agents are rewarded positively or negatively.
-Finally, lines 11-19 address monitoring length, visualization of agent performance, and resetting the environment if necessary.
+Algorithm 1 outlines the creation of the Predictive GraphRL Environment, a crucial component of the proposed GraphRL framework.
+This environment leverages the T-GCN, chosen for its effectiveness in capturing spatial-temporal dynamics essential for complex systems.
+The T-GCN’s ability to forecast future states, in addition to analyzing current and historical data, makes it invaluable for critical applications like health monitoring and traffic management, where early detection and timely response are key.
+A significant feature of this algorithm is its reward mechanism, which is instrumental in guiding the learning process.
+By awarding rewards based on the suitability of the agent’s actions, the environment ensures that the agent’s policy is aligned with the primary objectives of accurate forecasting and effective intervention.
+This design was motivated by the need for a proactive system, capable of not only forecasting but also informing real-time decision-making.
+The algorithm is meticulously structured to set up the observation space, action space, and reward policy based on predicted states.
+The initial lines (1-3) justify the use of the T-GCN model, especially for its applicability in dynamic and nonlinear data contexts like vital sign monitoring.
+The subsequent lines (4-5) are dedicated to initializing the environment, forming the basis for RL-driven decision-making.
+The reward policy, detailed in lines 6-10, aligns with standard RL practices, promoting actions that yield beneficial outcomes.
+Finally, lines 11-19 focus on continuous monitoring and adaptation, a critical aspect for applications that demand real-time responsiveness, such as in healthcare scenarios.
+This algorithm represents a significant step in advancing the field of RL, moving from passive observation to an active role in shaping decisions.
 IV-C GraphRL Agent
-In this study, the game learning agent Deep Q-Networks (DQN) algorithm is used.
-The algorithm was introduced by Google’s DeepMind for playing Atari game to play games by just observing the screen without any training or prior knowledge about those games.
-In this algorithm, the Q-Learning functions’ approximation is computed using the proposed T-GCN model and the learning agent gets rewarded based on the graph network prediction of the right action for the current state.
+In this study, the Deep Q-Networks (DQN) algorithm is used.
+The Deep Q-Networks (DQN) algorithm, developed by Google’s DeepMind, was initially designed for playing Atari games.
+This algorithm enabled the AI to learn game strategies directly from visual input, without requiring pre-programmed rules or prior game-specific training.
+In this algorithm, the Q-Learning functions are approximated using the proposed T-GCN model, and the learning agent is rewarded based on the graph network prediction of the right action for the current state.
 IV-C1 Q-Function Approximation
 T-GCN model used in this study to approximate the Q-Function for each action in the action space as shown in Fig.
-1.
+2.
 The model is configured with parameters such as the relu activation function, mean square error as loss function, and Adam optimiser.
@@
 The learning agent performs an action $a_{t}\epsilon A$ for a transition from state $s_{t}$ to $s_{t}^{{}^{\prime}}$ and achieves a reward $R$ for the action.
-In this transition process, the maximum of the Q-function in Equation 7 is calculated, and the discount of the calculated value uses a discount factor $\gamma$ to suppress future rewards and focus on immediate rewards.
+In this transition process, the maximum of the Q-function in Eq.
+7 is calculated, and the discount of the calculated value uses a discount factor $\gamma$ to suppress future rewards and focus on immediate rewards.
 The discounted future reward is added to the current reward to get the target value.
 The difference between the current prediction from the neural networks and the calculated target value provides a loss function.
-The loss function is a deviation of the predicted value from the target value and it can be estimated from Equation 8.
+The loss function is a deviation of the predicted value from the target value and it can be estimated from Eq.
+8.
 The square of the loss function allows for the punishment of the agent for a large loss value.
@@
 3:$memory\leftarrow append(s_{t},a_{t},R,s_{t+1})$
-4:if $np.random.rand<\epsilon$ then $\triangleright$ Exploration
+4:if $np.random.rand()<\epsilon$ then $\triangleright$ Exploration
 5:    $action\_value\leftarrow random(a_{t})$
@@
 12:    if not done then
-13:         $target\leftarrow R+\gamma(max(model.predict(s_{t+1})))$
+13:       $target\leftarrow R+\gamma\cdot\max(model.predict(s_{t+1}))$
 14:    end if
@@
 18:end for
-19:if $\epsilon$geq $\epsilon_{min}$ then
-20:     $\epsilon$= $\epsilon_{decay}$
+19:if $\epsilon\geq\epsilon_{min}$ then
+20:    $\epsilon\leftarrow\epsilon\cdot\epsilon_{decay}$
 21:end if
-The Learning Agent algorithm 2 encapsulates methods for function approximation, memorization, replay, exploration, and exploitation.
-In lines 1-2, hyperparameters required for the agent and T-GCN model for Q-function approximation are initialized.
-Memorization and replay, explained in line 3, involve storing T-GCN experience, including state, action, reward, and next_state, for later retraining of the model using the replay method.
-The algorithm (lines 4-8) predicts an action using either exploration or exploitation methods.
-Finally, lines 9-21 retrieve a batch of previous experiences from memory, process and retrain the neural network model based on the earlier defined hyperparameters.
+Algorithm 2 introduces the GraphRL Agent, presents the functionality of the GraphRL Agent within a complex action-state environment, utilizing T-GCN for Q-function approximation.
+This integration enables effective handling of spatial-temporal data complexities, enhancing decision-making.
+The algorithm’s design is founded on a strategic balance between exploration and exploitation, achieved via an epsilon-greedy strategy, crucial for adaptive learning and continual improvement in decision-making.
+It starts with initializing key parameters (Line 1), defining a T-GCN model for handling complex data structures (Line 2), and storing memories for experience replay (Line 3).
+The agent’s learning process involves iterative learning from minibatches of experiences, computing target Q-values, and adjusting its policy (Lines 8-21), with a dynamic adjustment of the exploration rate (Lines 16-18).
+The predict() function, pivotal in the exploitation phase, utilizes the T-GCN model’s predictions to guide actions, showcasing the algorithm’s advanced approach in navigating dynamic environments through a blend of exploration and strategic exploitation.
 IV-C4 Implementation Algorithm
@@
 1:Input:
-2:$\mathcal{C}={1,2,\dots,C}$: set of subjects
-3:$\mathcal{V}={1,2,\dots,V}$: set of vital signs
-4:$\mathcal{M}={1,2,\dots,M}$: number of episodes
+2:$\mathcal{C}=\{1,2,\ldots,C\}$: set of subjects
+3:$\mathcal{V}=\{1,2,\ldots,V\}$: set of vital signs
+4:=$\mathcal{M}=\{1,2,\ldots,M\}$: number of episodes
 5:Output: Rewards achieved by Agents in each episode.
@@
 9:    $state\leftarrow env.reset()$
-10:     $score=0$
+10:    $score\leftarrow 0$
 11:    for time in range(timesteps) do
@@
 22:end for
-The Algorithm  3 corresponds to the previous two algorithms,  1 and  2, and implements the proposed GraphRL monitoring framework.
-It takes as input a set of subjects with their vital signs and the number of training episodes for the agents.
-The output of the algorithm is the cumulative reward score achieved by the learning agent in each episode.
-The first two lines of the algorithm create instances of the Predictive GraphRL environment and GraphRL agent.
-Lines 3-17 consist of nested for loops with conditional statements to determine if an episode is complete.
-The outer loop iterates through each episode, resetting the environment to its initial state and the score to zero.
-The inner loop iterates through timesteps, which represent the current state’s time, and calls the methods defined in Algorithms  1 and  2 to predict the current state’s action, reward the agent for its predicted action, retrieve the next state, and store previous experiences.
-Finally, the replay method is called to update the neural network model with the stored experiences.
+Algorithm 3 serves as the comprehensive implementation of the GraphRL framework, intricately combining the Predictive GraphRL Environment (Algorithm 1) with the GraphRL Agent (Algorithm 2).
+This pivotal algorithm orchestrates the real-time interactions between the agent and the environment, thus forming the operational core of the framework.
+It outlines the simulation scope, input parameters (subjects $\mathcal{C}$, vital signs $\mathcal{V}$, and episodes $\mathcal{M}$), and the output in the form of cumulative rewards.
+The initialization phase prepares the environment env and the agent for interaction.
+The episodic loop, encompassing the agent’s action-response cycle, is vital for continuous learning and adaptation.
+Crucial to this process is the memorize() function, which stores experiences for later recall during experience replay, allowing the agent to learn from past actions and refine its decision-making strategy.
+This algorithm thus encapsulates the dynamic and iterative nature of the GraphRL framework, highlighting the importance of memory and experience in the realm of advanced reinforcement learning, and showcasing its functionality in complex, evolving environments.
 IV-D Bayesian Optimisation
 Bayesian optimisation is a global optimisation method that uses a probabilistic model to guide the search for the optimal solution.
-The model is updated as new data points are sampled and evaluated, allowing the algorithm to improve its predictions over time by fine-tuning the $L_{reg}$, $\lambda$ and minimising the loss function defined in Equation 6.
+The model is updated as new data points are sampled and evaluated, allowing the algorithm to improve its predictions over time by fine-tuning the $L_{reg}$, $\lambda$ and minimising the loss function defined in Eq.
+6.
 The basic idea behind Bayesian optimisation is to model the objective function, f(x), as a Gaussian process (GP).
 The GP model is used to predict the objective function value at any point x, given the observations of the function at other points.
-The prediction is given by the posterior distribution of the GP, which is a Gaussian distribution with mean and variance given by Equation 9.
+The prediction is given by the posterior distribution of the GP, which is a Gaussian distribution with mean and variance given by Eq.
+9.
 $\mu(x)$ $=k(x,X)^{T}(K+\sigma^{2}I)^{-1}y$ (9)
@@
 Common acquisition functions include the probability of improvement and the expected improvement.
-Given a set of observed points (X,y) and a Gaussian process prior, Bayesian optimisation seeks the point x* that minimises the loss function value, given by Equation 10.
+Given a set of observed points (X,y) and a Gaussian process prior, Bayesian optimisation seeks the point x* that minimises the loss function value, given by Eq.
+10.
 The optimisation process continues iteratively, sampling new points and updating the GP model until a stopping criterion is met.
@@
 V-A Datasets
-In this study, the proposed framework is evaluated with three publicly available datasets related to healthcare, traffic and weather.
-  • Wearable Stress and Affect Detection (WESAD) dataset [32]: This is a publicly available dataset that contains physiological and self-report data collected from 17 participants during various stress and affective states.The dataset contains data collected from various wearable sensors, including a chest-mounted electrocardiogram (ECG) sensor, a wrist-mounted photoplethysmogram (PPG) sensor, and an accelerometer.
-  • Los Angeles (LA) traffic dataset [33]: This is a publicly available dataset that contains data on traffic flow in the Los Angeles metropolitan area.
-The dataset was collected by the Los Angeles Department of Transportation (LADOT) and contains information on traffic counts, traffic speeds, and travel times for a variety of purposes, including traffic prediction, traffic flow analysis, and transportation planning.
-  • Large-Scale Traffic and Weather Events (LSTW) dataset [34]: It contains information on traffic conditions and extreme weather events in the United States.
-In terms of weather events, the dataset has several types including rain, snow, storm, cold weather event, etc.
+The GraphRL framework’s testing with datasets from healthcare, traffic, and weather domains was a deliberate strategy to assess its versatility and robustness in handling diverse time-series data.
+The choice of these varied domains was intended to demonstrate the framework’s adaptability and efficacy in different contexts.
+Each domain poses unique challenges: healthcare data’s complexity and sensitivity, traffic data’s dynamic patterns requiring real-time analysis, and weather data’s intricate interplay of environmental factors.
+Successfully navigating these distinct datasets underscores the framework’s capability for widespread real-world application.
+Additionally, using datasets from different fields facilitates a thorough evaluation of the framework, ensuring its versatility and effectiveness across various problem types and data structures.
+This comprehensive approach is vital for a tool designed for extensive applications in data analysis and prediction.
+Three datasets utilized for evaluating the GraphRL framework as shown in Fig.
+I, each from a different domain: healthcare, traffic, and weather.
+In healthcare, the WESAD dataset, containing electrocardiogram (ECG) and photoplethysmogram (PPG) data from 17 participants, offers a rich source of biometric time-series data for pattern recognition analysis.
+The Los Angeles (LA) Traffic dataset, sourced from the Los Angeles Department of Transportation (LADOT), provides real-time urban traffic data like traffic counts and speeds, while the Large-Scale Traffic and Weather Events (LSTW) dataset, with data across the United States, uniquely combines traffic conditions and weather events, posing a multifaceted challenge for the framework.
+TABLE I: Datasets
+Dataset Domain Key Features Statistics Suitability
+WESAD
+[26]
+Healthcare
+Physiological data
+(ECG, PPG),
+motion data
+17 participants
+Rich biometric time-series
+data, ideal for testing pattern
+recognition in health-related
+data.
+LA
+Traffic
+[27]
+Traffic
+Traffic counts,
+speeds,
+travel times
+Data from
+LADOT
+Real-time urban traffic data,
+useful for analyzing and
+predicting dynamic traffic flows.
+LSTW
+[28]
+Weather
+Traffic conditions,
+weather events
+Data across
+the United
+States
+Combines traffic and weather
+data, challenging the framework
+to handle complex, multifactorial
+scenarios.
 V-B Baseline Models
-  • In Li et al.’s work [35], the authors suggested the ELMA method for multi-agent activity forecasting through graph neural networks that incorporate spatiotemporal data using simultaneous energy-based learning.
-  • Ma et al. [36] proposed a technique for predicting multi-agent driving behaviour in different scenarios based on self-supervised domain knowledge.
-  • In Jiang et al.’s research [37], they conducted a study on internet traffic prediction with distributed multi-agent learning and utilized both LSTM and gated recurrent unit (GRU).
-Their GRU-based distributed multi-agent learning scheme outperformed LSTM.
+In our study, we selected three baseline models for comparison, each epitomizing state-of-the-art approaches in multi-agent forecasting, graph neural networks, and traffic prediction.
+These models were chosen based on their innovative methodologies and proven effectiveness in areas closely aligned with our research objectives.
+  • ELMA Method [29]: Developed by Li et al.
+[29], the ELMA method utilizes graph neural networks for forecasting multi-agent activities, particularly adept at handling spatiotemporal data.
+Its novelty lies in the use of energy-based learning, making it an excellent benchmark against our framework, which similarly leverages graph-based techniques in complex environments.
+  • Self-Supervised Technique [30]: This technique is pioneering in self-supervised learning for predicting multi-agent driving behavior.
+Its relevance to our study comes from its focus on behavior prediction in diverse scenarios, using self-supervised domain knowledge—an advanced trend in multi-agent learning.
+  • Internet Traffic Prediction Study [31]: It involves internet traffic prediction using distributed multi-agent learning, employing LSTM and GRU models.
+GRU’s superior performance in their study provides a valuable point of comparison for our research, which focuses on sophisticated learning techniques in traffic prediction.
+Each of these models represents a significant stride in their respective fields.
+Their selection for comparison in our study is justified by their alignment with our research goals and their benchmark status in handling complex, dynamic datasets.
+By comparing our GraphRL framework against these models, we aim to demonstrate our approach’s novelty and effectiveness in diverse real-world applications.
 V-C Evaluation Metrics
-Mean Absolute Error (MAE) is a commonly used regression metric that determines the average magnitude of errors between the predicted and actual values in a given dataset.
-It is calculated by taking the average of the absolute differences between the predicted and actual values, which results in a single value.
-Another popular regression metric is the Root Mean Squared Error (RMSE), which measures the average magnitude of the differences between the predicted and actual values.
-RMSE is calculated as the square root of the mean of the squared differences between the predicted and actual values.
-Mean Absolute Percentage Error (MAPE) is another regression metric that calculates the average absolute percentage error between the predicted and actual values.
-It is calculated as the average of the absolute differences between the predicted and actual values, expressed as a percentage of the actual values.
-Cumulative Rewards, on the other hand, is a performance metric used in reinforcement learning that quantifies the total rewards obtained by an agent within a specific time frame or number of actions.
-It is computed by summing up all the rewards that the agent received during the given period or a number of actions.
-The experiments conducted in this study employed Python version 3.7.6 and utilized several packages, including TensorFlow, Keras, Open AI Gym, and stable_baselines3.
-TABLE I: Performance of the proposed framework in health forecasting
+Mean Absolute Error (MAE) is a widely-utilized regression metric that gauges the average magnitude of errors between predicted and actual values in a dataset.
+It is calculated by averaging the absolute differences between these values, yielding a singular metric.
+Root Mean Squared Error (RMSE) is another prominent regression metric, assessing the average magnitude of differences between predicted and actual values.
+RMSE is computed as the square root of the mean of these squared differences.
+Mean Absolute Percentage Error (MAPE) represents yet another regression metric, quantifying the average absolute percentage error between predicted and actual values.
+It is derived by averaging the absolute differences between these values, expressed as a percentage of the actual values.
+Conversely, Cumulative Rewards is a performance metric specific to reinforcement learning.
+It measures the total rewards an agent accumulates over a specified timeframe or across a set number of actions, calculated by summing all rewards received during this period.
+In the context of the experiments conducted for this study, Python version 3.7.6 served as the programming environment, with the deployment of several packages including TensorFlow, Keras, OpenAI Gym, and stable_baselines3.
+TABLE II: Performance of the proposed framework in health forecasting
 15 Min 30 Min 45 Min 60 Min
@@
 The proposed T-GCN in the predictive GraphRL environment performs better than the other baseline frameworks ELMA, GRU, and GRU-Based Multi-Agent as shown in Tab.
-I.
+II.
 It achieves the lowest MAE, MAPE and RMSE values in all the time intervals.
-TABLE II: Performance of the proposed framework in traffic forecasting
+TABLE III: Performance of the proposed framework in traffic forecasting
 15 Min 30 Min 45 Min 60 Min
@@
 As shown in Tab.
-II, T-GCN outperforms the other models for all the forecasting intervals with the lowest MAE, MAPE, and RMSE values.
+III, T-GCN outperforms the other models for all the forecasting intervals with the lowest MAE, MAPE, and RMSE values.
 The second-best performer is the GRU-Based Multi-Agent model, followed by GRU and ELMA.
-TABLE III: Performance of the proposed framework in weather forecasting
+TABLE IV: Performance of the proposed framework in weather forecasting
 15 Min 30 Min 45 Min 60 Min
@@
 The results show that the T-GCN model had the lowest MAE, MAPE, and RMSE values for all forecasting intervals, indicating its superior forecasting performance compared to the other models as shown in Tab.
-III.
+IV.
 VI-B GraphRL Agent Performance
-TABLE IV: Proposed GraphRL Performance
+TABLE V: Proposed GraphRL Performance
 AI Agents WESAD
@@
 The proposed RL agent was enabled with T-GCN and its performance is compared with other traditional RL agents as shown in Tab.
-IV.
+V.
 The table provides a comparison of different AI agents and their performance on three different datasets: WESAD, LAM Traffic Forecasting, and US Weather Forecasting.
@@
 The GraphRL agent has outperformed other RL agents in all three predictive and monitoring applications.
-[ xlabel=Episodes, ylabel=Score, legend pos=south east, legend style=font=, yticklabel style=/pgf/number format/fixed, ymin=-20000, ymax=65000, width=height=6cm, xtick=1,2,3,4,5,6,7,8,9,10, xticklabels=1,2,3,4,5,6,7,8,9,10]
-[color=blue, mark=*] coordinates (1, 38925) (2, 43595) (3, 46875) (4, 57280) (5, 32245) (6, 485) (7, 48035) (8, 48790) (9, 42885) (10, 54470) ;
-[color=red, mark=square*] coordinates (1, -8810) (2, -8100) (3, -11845) (4, 31960) (5, 46295) (6, 32560) (7, 53640) (8, 36195) (9, 36480) (10, 46425) ;
-[color=green, mark=triangle*] coordinates (1, 23780) (2, -6765) (3, 33480) (4, 43635) (5, 24925) (6, 2670) (7, 8990) (8, 53145) (9, 58530) (10, 41405) ;
-Health, Traffic, Weather
+[…]
 Figure 3: GraphRL Agent Rewards Distribution
@@
 The breakdown of the proposed agent’s score in each episode of the three applications is presented and compared in Fig.
-2.
+3.
 The agent’s performance on the WESAD dataset is relatively consistent, with the scores fluctuating between 32245 and 57280.
@@
 VII Bayesian Optimisation Results
-[ xlabel=Episode, ylabel=Rewards, legend pos=north west, legend style=font=, legend cell align=left, ymajorgrids=true, grid style=dashed, ymin=-15000, ymax=60000, scaled y ticks = false, y tick label style=/pgf/number format/sci, width=height=6cm]
-[…]
-[…]
-[…]
-[…]
 […]
 Figure 4: Bayesian optimisation of $\alpha$ for GraphRL Agent
-[ xlabel=Episode, ylabel=Rewards, legend pos=north west, legend style=font=, legend cell align=left, ymajorgrids=true, grid style=dashed, ymin=-15000, ymax=60000, scaled y ticks = false, y tick label style=/pgf/number format/sci, width=height=6cm]
-[color=red,mark=x] coordinates (1, -944) (2, -667) (3, 919) (4, 4009) (5, 7860) (6, 14303) (7, 21405) (8, 28799) (9, 38095) (10, 48823) ;
-[color=blue,mark=x] coordinates (1, -588) (2, 2080) (3, 3122) (4, 6573) (5, 9759) (6, 16085) (7, 25136) (8, 34720) (9, 46093) (10, 57388) ;
-[color=green,mark=x] coordinates (1, 909) (2, -258) (3, 1398) (4, 5207) (5, 10434) (6, 17728) (7, 25584) (8, 34898) (9, 32468) (10, 42637) ;
-[color=orange,mark=x] coordinates (1, -84) (2, 272) (3, 2355) (4, 5792) (5, 11162) (6, 17760) (7, 23002) (8, 31702) (9, 40771) (10, 52126) ;
 […]
@@
 The results of Bayesian optimisation for the proposed GraphRL agent using different values of the learning rate parameter, $\alpha$, during different episodes are shown in Fig.
-3.
+4.
 The values in the y-axis of the line chart represent the scores or rewards obtained by the agent during each episode.
@@
 VIII Conclusion
-In conclusion, this study proposed a novel GraphRL framework based on a Temporal Graphical Convolutional Network (T-GCN) to forecast the future states of a dynamic RL environment and overcome the limitations of deep learning models in time-series prediction tasks.
-The proposed framework was fine-tuned with the Bayesian optimisation technique and was applied to three datasets: WESAD, LAM Traffic Forecasting, and US Weather Forecasting.
-The results showed that the proposed GraphRL agent outperformed other traditional RL agents in all three predictive and monitoring applications.
-The agent’s performance on the WESAD dataset was relatively consistent, while on the LAM Traffic Forecasting and US Weather Forecasting datasets, the agent’s performance was relatively inconsistent.
-The inconsistency of the scores was due to the exploration rate.
-The Bayesian optimisation process was able to fine-tune the $L_{reg}$, $\lambda$ and minimise the loss function to improve the performance of the GraphRL agent.
-The proposed GraphRL framework can be applied to a wide range of applications and can be further enhanced by incorporating other forms of information such as node and edge attributes for more accurate predictions.
-Ethical Statement
-There are no ethical issues.
+The GraphRL framework, introduced in this study, embodies an innovative amalgamation of T-GCN and RL.
+It is specifically engineered to augment the prediction of future states in dynamic environments.
+Rigorous evaluations, utilizing an array of datasets such as WESAD, LA Traffic Forecasting, and US Weather Forecasting, have substantiated the framework’s enhanced performance compared to conventional RL agents.
+Nonetheless, it is imperative to acknowledge that the efficacy of GraphRL is significantly contingent upon the caliber of the input data and necessitates substantial computational resources.
+The framework’s reliance on data of high quality and structure constitutes a considerable limitation, with its accuracy and effectiveness being closely bound to the data’s integrity.
+Additionally, the computational requisites, predominantly due to the T-GCN model integration, present challenges in scalability and broader applicability.
+Future enhancements of the GraphRL framework will be directed towards surmounting these constraints and broadening its functional scope.
+Prospective developments entail the incorporation of spatial data processing, aimed at bolstering the framework’s analytical prowess, particularly in processing data with spatial or geographical dimensions.
+Investigating a spectrum of graph-based models could yield insights for enhancing both the efficiency and efficacy of the framework.
+Furthermore, the exploration of real-time adaptive learning strategies presents a promising avenue for subsequent research.
+Such advancements are anticipated to enable the framework to dynamically adapt to evolving data patterns and environmental shifts.
+In summation, the GraphRL framework signifies a substantial advancement in the domain of time-series prediction and monitoring.
+Its adeptness in managing complex temporal data surpasses traditional RL methodologies, heralding innovative applications in sectors such as healthcare, traffic management, and environmental forecasting.
+As the framework undergoes continued refinement and evolution, it is positioned to emerge as an instrumental component in the progression of predictive analytics and intelligent monitoring systems, with extensive applicability across diverse fields.